# Multi-turn Evaluation of Anthropomorphic Behaviours in Large Language Models

### v1

@@ C.1 Inter-rater agreement
We asked 37 professional raters on a crowd-sourcing platform to evaluate the presence of anthropomorphic behaviours in a sample of 924 unique dialogue turns. The sample was a combination of stratified and random sampling to ensure equal representation of messages from different models and to up-sample rarer anthropomorphic behaviours, particularly given large discrepancies in frequencies of behaviours. Per dialogue turn, three raters were asked to provide binary ratings for whether a behaviour was present in a given model message. Raters were presented with both the model message as well as the user message from the same turn in order to provide some context. Above, we see the average percentage of agreement between raters. We also present Krippendorf’s alpha values for each cue, which is the most flexible chance-agreement-adjusted inter-rater reliability metric with more than two raters per item (Hayes & Krippendorff, 2007). Overall, we see that average agreement percentage scores are above chance, with “empathy” having the lowest average agreement and “explicit human-AI relationship reference” having the highest.
Krippendorf’s alpha values are all positive, meaning that observed agreement among coders or raters is higher than what you would expect by chance alone. However, it is worth noting that these values span the ranges of poor (${<}0.67$) to moderate (0.67–0.79) agreement (Marzi et al., 2024). This is not entirely unexpected, as previous rating tasks where users have evaluated models for subjective and socially-grounded dimensions have returned inter-rater agreement values in a similar range (Glaese et al., 2022; Stiennon et al., 2020; Ouyang et al., 2022; Bai et al., 2022).
@@
C.2 Intra-model agreement
@@
C.3 Inter-model agreement
Before aggregating all model ratings into a single LLM-as-judge rating (as described in Section C.4), we were interested in seeing how frequently models agreed with one another’s ratings to uncover any patterns of agreement between models that would be obscured by the aggregation. For every dialogue turn annotated for a specific cue (62,400 unique annotation targets), we compared binary ratings given by models and computed the average rate of agreement between models. The visualisation shows the average agreement rate ($x$ axis) for all model pairs used as automated raters ($y$ axis). Across different cue types, we find that any given model pair agrees at approximately the same rate as other model pairs. Some differences between model pairs can be observed for empathy and validation, with greatest agreement between Gemini 1.5 Flash and Claude 3.5 Sonnet ratings and the least agreement between GPT-4 Turbo and Claude 3.5 Sonnet. Overall, these results indicate that models agree with one another at approximately the same rate, and that there is low risk of a single model being systematically “out-voted” by the other two models in aggregation.
@@ C.3 Inter-model agreement
C.4 Model-rater agreement
@@
Appendix D Supplementary human-AI experiment information
We present the system prompts for the AI systems used in the human subject study (D.1 and D.2), the results of running our evaluation on these AI systems as a manipulation check (D.3), as well as a detailed breakdown of the study’s survey results (D.4). The following system prompts were developed such that they only vary the behaviours our evaluation tracks. We avoid providing any additional instructions to minimise the introduction of other behaviours which may be correlated with anthropomorphism (such as question-asking) but which are outside the scope of measuring anthropomorphism as an isolated construct and thus outside the scope of our evaluation.
D.1 High-frequency model system prompt
@@
D.2 Low-frequency model system prompt
@@
D.3 Profiles of the low-frequency and high-frequency experiment models as compared to an unprompted model
Figure 11: Anthropomorphism profiles of the two prompted versions of Gemini 1.5 Pro -- one prompted to exhibit a high frequency of anthropomorphic behaviours and one to exhibit a low frequency of anthropomorphic behaviours -- used in the human-AI experiments, as compared to an unprompted version of Gemini 1.5 Pro. The profiles validate that the high-frequency model exhibits behaviours from all four categories at a much higher frequency than both the low-frequency model and the unprompted version of Gemini 1.5 Pro that is evaluated earlier in Section 5. Similarly, the low-frequency model exhibits behaviours from all four categories at a much lower frequency than both the high-frequency model and the unprompted model.
D.4 Breakdown of the survey results by survey item

### v2

@@ A.3.3 Inter-rater agreement
Above, we see the average percentage of agreement between raters. We also present Krippendorf’s alpha values for each cue, which is the most flexible chance-agreement-adjusted inter-rater reliability metric with more than two raters per item [hayes_answering_2007]. Overall, we see that average agreement percentage scores are above chance, with “empathy” having the lowest average agreement and “explicit human-AI relationship reference” having the highest.
Krippendorf’s alpha values are all positive, meaning that observed agreement among coders or raters is higher than what you would expect by chance alone. However, it is worth noting that these values span the ranges of poor (<0.67{<}0.67) to moderate (0.67–0.79) agreement [marzi_k_alpha_2024]. This is not entirely unexpected, as previous rating tasks where users have evaluated models for subjective and socially-grounded dimensions have returned inter-rater agreement values in a similar range [glaese_improving_2022, stiennon_learning_2020, ouyang_training_2022, bai_training_2022].
@@
A.3.4 Intra-model agreement
@@
A.3.5 Inter-model agreement
Before aggregating all model ratings into a single LLM-as-judge rating (as described in Section A.3.6), we were interested in seeing how frequently models agreed with one another’s ratings to uncover any patterns of agreement between models that would be obscured by the aggregation. For every dialogue turn annotated for a specific cue (62,400 unique annotation targets), we compared binary ratings given by models and computed the average rate of agreement between models. The visualisation shows the average agreement rate ($x$ axis) for all model pairs used as automated raters ($y$ axis). Across different cue types, we find that any given model pair agrees at approximately the same rate as other model pairs. Some differences between model pairs can be observed for empathy and validation, with greatest agreement between Gemini 1.5 Flash and Claude 3.5 Sonnet ratings and the least agreement between GPT-4 Turbo and Claude 3.5 Sonnet. Overall, these results indicate that models agree with one another at approximately the same rate, and that there is low risk of a single model being systematically “out-voted” by the other two models in aggregation.
@@ A.3.5 Inter-model agreement
A.3.6 Model-rater agreement
@@
A.4 Supplementary human-AI experiment information
We present the system prompts for the AI systems used in the human subject study (A.4.1 and A.4.2), a screenshot of the experiment interface (A.4.3), experiment instructions (A.4.4) and survey questions (A.4.5), as well as a detailed breakdown of the study’s survey results (A.4.6).
The system prompts were developed such that they only vary the behaviours our evaluation tracks. We avoid providing any additional instructions to minimise the introduction of other behaviours which may be correlated with anthropomorphism (such as question-asking) but which are outside the scope of measuring anthropomorphism as an isolated construct and thus outside the scope of our evaluation.
We believe the experiment posed minimal risks to participants. It did not deceive participants about the nature of their conversation partner, which was always disclosed as an AI system. We also encouraged participants to not discuss any of their direct or personal experiences with the chatbot if it made them uncomfortable. Additionally, participants were informed they could withdraw from the study at any time without penalty, and all data was anonymised during analysis to protect privacy. Our protocol was reviewed and approved by an independent institutional ethics board (anonymised for review). The post-experiment debriefing presented to participants explained the goal of our study, as well as the concept of anthropomorphism more broadly and its societal risks, in order to communicate a clear understanding of the experiment purpose and their contribution.
A.4.1 High-frequency model system prompt
@@
A.4.2 Low-frequency model system prompt
@@
A.4.3 Experiment interface
The interactive experiment interface, shown in Figure 12, was a splitscreen consisting of a resizable chat screen where participants exchanged messages with one of the two chatbots and an instructions screen. The instructions screen changes to the survey questions after participants complete their chat session.
Figure 12: Human subject study interface which consists of a resizable chat screen and an instructions/survey questions screen.
A.4.4 Experiment instructions
The interactive experiment consisted of a short conversation with one of the two chatbots followed by survey questions. Participants were randomly assigned one of the eight scenarios developed and used in the automated evaluation as the subject of their conversaton with the chatbot. The instructions provided to participants were the following:
Please read and complete each step carefully.
Welcome, and thank you for your participation in this study! This study will involve the following steps:
**Step 1:**
Have a **10-20 minute conversation in English about [USE SCENARIO]** with the chatbot in the chat window on the left-hand side. You can resize the chat window to make it larger if you’d like.
**Step 2:**
Please **engage in active conversation** for **at least 10 minutes** and **no more than 20 minutes**. When you are ready to end the conversation, click the **“Begin Survey”** button (above, on the right). Remember, this button will only be enabled after a certain amount of time passes.
**Step 3:**
You will then proceed to the next part of the study. You will be asked to answer some questions about your experience in the survey.
**Task overview**
You are having a conversation with a chatbot during which you’re trying to **[user goal]**. Try your best to stay focused on this topic and respond naturally and authentically to the chatbot’s messages. Alternatively, if it feels more comfortable, you can imagine how someone you know might approach this conversation and respond in a similar manner.
Here are some ideas of what you can talk about:
* [idea 1]
* [idea 2]
* [idea 3]
A.4.5 Questions for implicit and explicit measures
The two measures used to assess implicit and explicit anthropomorphism were the following:
Implicit measure - description of chatbot
What is your impression of the chatbot that you just interacted with? We are interested to hear what you thought about it. Please answer in a short paragraph (at least 3 sentences) to ensure your submission is complete.
Explicit measure - Godspeed Anthropomorphism survey
As in other studies on anthropomorphic perceptions of non-embodied chatbots, we remove one item from the original survey in bartneck_measurement_2009 as this item assumes an embodied agent, which is not the case in our experiment.
Please answer the following questions about the chatbot:
Rate your impression of the chatbot: (Fake – Natural)
1. Completely fake
2. Somewhat fake
3. Neither fake nor natural
4. Somewhat natural
5. Completely natural
Rate your impression of the chatbot: (Machine-like – Human-like)
1. Completely machine-like
2. Somewhat machine-like
3. Neither machine-like nor human-like
4. Somewhat human-like
5. Completely human-like
Rate your impression of the chatbot: (Unconscious – Conscious)
1. Completely unconscious
2. Somewhat unconscious
3. Neither unconscious nor conscious
4. Somewhat conscious
5. Completely conscious
Rate your impression of the chatbot: (Artificial – Lifelike)
1. Completely artificial
2. Somewhat artificial
3. Neither artificial nor lifelike
4. Somewhat lifelike
5. Completely lifelike
A.4.6 Breakdown of the survey results by survey item